OpenStudy (anonymous):

When finding a derivative using the definition of a derivative, how do you know when the use the formula lim as y approaches x f(y)-f(x)/(y-x) vs. lim as h approaches 0 f(x+h)-f(x)/(h) ?

OpenStudy (freckles):

you can use either one

OpenStudy (freckles):

it is which ever one you like most

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

Both formulas are actually the same. Try replacing \(h\) with \(y-x\) to see what happens.

OpenStudy (freckles):

example: \[f(x)=x^3 \\ f'(x)=\lim_{y \rightarrow x}\frac{f(y)-f(x)}{y-x}=\lim_{y \rightarrow x} \frac{y^3-x^3}{y-x}=\lim_{y \rightarrow x} \frac{(y-x)(y^2+xy+x^2)}{y-x} \\ f'(x)=\lim_{y \rightarrow x} (y^2+xy+x^2)=x^2+x(x)+x^2=3x^2 \\ \text{ or } f'(x)=\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}=\lim_{h \rightarrow 0}\frac{(x+h)^3-x^3}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{x^3+3x^2h+3xh^2+h^3-x^3}{h}=\lim_{h \rightarrow 0} \frac{3x^2h+3xh^2+h^3}{h} \\ f'(x)=\lim_{h \rightarrow 0}(3x^2+3xh+h^2)=3x^2+3x(0)+0^2=3x^2\]

OpenStudy (freckles):

the first method involved factoring and the second one involved multiplying

OpenStudy (freckles):

it just depends on in that case if you preferred multiplication over the factoring some people forget how to factor a difference of cubes

OpenStudy (freckles):

thought I hate multiplying sometimes

OpenStudy (freckles):

but I didn't really multiplied I sorta remember that line in the pascal's triangle :p

OpenStudy (anonymous):

lol okay. thank you for the clear example. i just tried one myself using both methods and got the same answer.

OpenStudy (freckles):

sith mentioned proving they are equivalent

OpenStudy (freckles):

if you want to do that

OpenStudy (freckles):

or try that

OpenStudy (freckles):

\[(x-h)^3-x^3=([x+h]-[x])([x+h]^2+[x+h][x]+[x]^2) \\ (x+h)^3-x^3=h((x+h)^2+x(x+h)+x^2)\] and another option for that second method like we could have factored the numerator too